# Quantifying community evolution in temporal networks

### v2

@@ Abstract
When we detect communities in temporal networks it is important to ask questions about how they change in time. Normalised Mutual Information (NMI) has been used to measure the similarity of communities when the nodes on a network do not change. We propose two extensions namely Union-Normalised Mutual Information (UNMI) and Intersection-Normalised Mutual Information (INMI). UNMI and INMI evaluate the similarity of community structure under the condition of node variation. Experiments show that these methods are effective in dealing with temporal networks with the changes in the set of nodes, and can capture the dynamic evolution of community structure in both synthetic and real temporal networks. This study not only provides a new similarity measurement method for network analysis but also helps to deepen the understanding of community change in complex temporal networks.
I INTRODUCTION
@@
An important problem in network science is understanding how communities form in networks. Communities within these networks do not remain static: they expand, merge, or dissolve, reflecting the underlying changes in relationships and interactions [38, 16, 8, 12]. We might hypothesise that some networks have relatively stable and long-lasting relationships but other networks might form only transient communities, but how can this be measured? A community here is a partition of a network into (usually) distinct sets of nodes with connections more common within communities. A common measure used on such a partition is modularity [25, 26, 13], though it has attracted some criticism [30]. When we consider a network that changes in time we can choose to analyse the communities only within a time window and ask how much communities change between such windows. This requires us to rigorously compare two different partitions of a network and to give a measure of how similar those partitions are. This measure must allow for the fact that nodes can leave or join the network.
@@
A well-known method for comparing two such partitions is normalised mutual information (NMI) but this measure assumes that both networks have the same set of nodes. The primary challenge addressed in this research is the accurate comparison of community structures in networks where the set of nodes varies over time. This study proposes variants of traditional NMI: Union-Normalised Mutual Information (UNMI) and Intersection-Normalised Mutual Information (INMI) which compares two networks with different node sets either by considering the union or intersection of the node sets. We show by using artificial models that these measure different aspects of how communities change over time and demonstrate their utility on artificial networks with known characteristics. Following this we use real temporal network data sets which we split into different time windows. We demonstrate how graph statistics can be used to pick an appropriate time window for the task. We show that the methods well recover the known properties of artificial data. We show that on real data the length of stability of communities found is short compared with our artificial models but changes for different data sets.
I-A Related work
@@
The focus of this paper is on how communities change rather than specifically on how to partition a network into communities. Here we begin with a basic assumption, commonly made, that a community detection algorithm partitions all nodes in a network into a set of non-overlapping communities. Community detection is far from a solved problem. The problem can be divided into inferential methods and descriptive methods [30]. Inferential methods attempt to fit parameters in some predetermined model that generates communities. Descriptive methods attempt to partition the network to maximise some measure of how good that partition is at creating communities. Modularity is commonly used for the latter as it attempts to measure the proportion of links that start and end in the same community rather than that which would be expected by a null model. In this paper, the measures we use do not rely on whether descriptive or inferential techniques are available. We use the Louvain algorithm to detect communities for our investigation into real-world data because it is well-known, relatively quick to run and simple to understand. However, the core contribution of the UNMI and INMI and their use to measure how fast communities change in time would not change if a different community detection method were used [33, 36]. For the work on artificial networks, we simply assume that the communities exist (community membership is generated directly from the model) and that some algorithm could in principle detect them.
@@
The dynamic characteristics of dynamic communities play an important role in the study of the evolution of complex systems, which provides a bridge for the interaction between micro and macro structures in complex systems. The paper [11] discusses the stability of the community structure in networks and proposes a stability measurement method based on the Markov process. This method reveals the community structure characteristics of the network at different resolutions through time-scale analysis. Paper [4] highlights the relationship between nodes’ dynamics with the formation and evolution of community structures in dynamic environments. These works reveal the dynamic characteristics of dynamic community structure to some extent. Based on existing studies, we improved existing techniques and compared the similarity of community structure of temporal networks at different time points on a one-to-one basis. A direct approach to comparing the similarity of the network at the mesoscale level is comparing the similarity of the partition of nodes. Normalised Mutual Information (NMI) [9] is a widely used method to measure the similarity of partitions, by comparing the similarity of two partitions by measuring information gain. Recent studies have pointed out that the normalised parameters of NMI are biased, and its variants [20, 40, 1] and asymmetric normalise technique [20] are proposed to solve the problem. The Rand Index and its variants [34, 19] and the Wallace coefficient [41, 37] are less commonly used methods which assess the difference between partitions. Like NMI they require the same node set for both partitions compared. Partition edit distance (PED) [2] is an extension of graph edit distance that defines operands to manipulate the partition of nodes. However, it relies on one-to-one matching between communities, and the high computational overhead also limits the application of this method on large-scale networks.
In addition to the methods discussed above, some candidate techniques could compare the similarity (or distance) of the dynamic community structure of temporal networks. Examples include graph-based spectral distances [39, 42, 21] or graph-based embedding techniques [14, 32, 15, 17]. However, these methods are not directly designed for the task and are less interpretable and computationally efficient than the methods discussed above.
II METHODS
@@
Comparing network partitions is a fundamental task in network science and community detection. Normalized Mutual Information (NMI) is one of the most widely used metrics due to its ability to effectively quantify the similarity between partitions. However, traditional NMI assumes that the partitions being compared share an identical set of nodes, which often is not the case in real-world scenarios where networks may have differing node sets due to growth, decay, or entirely different origins. This limitation becomes evident when attempting to compare partitions of networks with varying node compositions, rendering traditional NMI inadequate for such analyses. Recognizing this gap, we propose an approach that explicitly compares network partitions with different node sets by introducing two novel metrics: Union-NMI (UNMI) and Intersection-NMI (INMI). This approach is motivated by the practical need to assess the similarity between partitions in networks where node addition or deletion is common, such as in temporal networks or comparative studies across different systems.
@@
In our framework, UNMI considers the union of the node sets from the partitions being compared, effectively incorporating all nodes present in either partition. This allows for a comprehensive evaluation of similarity that accounts for the entirety of both networks. Conversely, INMI focuses on the intersection of the node sets, assessing similarity based solely on the nodes common to both partitions.
II-A Temporal networks model
@@
We adopt the snapshot model [18] of a sequential network to study the dynamic evolution of networks in time dimension. The model works by dividing the entire temporal network into a series of discrete time segments, each corresponding to a snapshot of the temporal network called a snapshot. This method allows us to observe changes in the structure of the network at different points in time and analyse the dynamic behaviour of nodes and edges. Formally, for a temporal network $G_{T}=(V,E_{T},t_{0},t_{\max})$ that is observed between $t_{0}$ and $t_{\max}$, it can be represented by a series of snapshots where each slice contains the temporal events occur at the time window $T_{i}\in[t,t+\tau)$, denoted by $G_{T}=(G_{1},G_{2},\dots,G_{n})$, where $\tau$ is the size of the time window, each snapshot $G_{i}=(V_{i},E_{i})$ represents the subgraph in time window $T_{i}$. $V_{i}\subseteq V$ is the nodes activate in time window $T_{i}$ and $E_{i}\subseteq\{(u,v)\in E_{T}|t\in T_{i}\}$ is the events happen in time window $T_{i}$. In addition, we used overlapping time windows to build snapshots from a temporal network, which allowed us to build “smooth” snapshot sequences. This means that the network structure of two contiguous snapshots does not mutate.
Choosing the right window size is critical to building a snapshot model of a temporal network. The small window size makes the network too sparse, thus ignoring the potential community structure; A large window size will mask the dynamic nature of the network. Although several methods for extracting timescale in dynamic datasets already exist [10, 29, 5, 31], given that our problem is choosing a reasonable time window to build a network with meaningful community structure, we analyze the modularity of the network under different windows and the proportion of the largest connected component to select the appropriate time window. Intuitively, in a snapshot model constructed with a reasonable time window, the network should not be too sparse in each time slice, and the modularity of each time slice should be large, indicating that there is a meaningful community structure in the network at this time. We measured the statistical properties of a temporal network in different time windows from multiple indicators: the modularity, the proportion of LCC, the edge-node ratio of slices in different time windows, and the Z-score of modularity compared with null models respectively. We describe this method in detail in supporting information and list the experimental data and reasons why we chose the time window for each data set.
II-B Z-score calculation
Choosing the right window size is critical to building a snapshot model of a dynamic network. The small window size makes the network too sparse, thus ignoring the potential community structure; A large window size will mask the dynamic nature of the network. Although several methods for extracting timescale in dynamic datasets already exist [10, 29, 5, 31], given that our problem is choosing a reasonable time window to build a network with meaningful community structure, we analyse the modularity of the network under different windows and the proportion of the largest connected component to select the appropriate time window. Intuitively, in a snapshot model constructed with a reasonable time window, the modularity of each time slice should be large, indicating that there is a meaningful community structure in the network at this time.
Our observations show that using modularity alone does not fairly compare the degree to which nodes form communities between two different networks. What we found is that in more sparse networks, modularity is always higher. The reasons for this are varied, but one possible reason is that the lower density divides the network into natural submodules, and these natural submodules contribute to higher modularity, yet they cannot be interpreted as meaningful community structures. Recalling the definition of modularity of a partition of the node set of a network:
$Q=\frac{1}{2m}\sum_{ij}\left(A_{ij}-\gamma\frac{k_{i}k_{j}}{2m}\right)\delta\left(c_{i},c_{j}\right),$
where $m$ is the number of edges (or sum of all edge weights), $A$ is the adjacency matrix of $G$, $k_{i}$ is the (weighted) degree of node $i$, $\gamma$ is the resolution parameter, and $\delta(c_{i},c_{j})$ is 1 if node $i$ and node $j$ are in the same community else 0. Note that the density of the network affects the term $\gamma\frac{k_{i}k_{j}}{2m}$. As a result, higher results are often obtained when modularity is calculated on a network with lower density. Therefore, when the density of the two networks involved in the comparison is different, it is unfair to compare modularity simply. Generally, a larger window size tends to produce a denser network. Therefore, we use null models with the same degree sequence for comparison, thus eliminating the influence of network density on modularity. In the methods section, we went into more detail about how to build a null model by configuration model [25] and calculate the Z-score.
To evaluate the significance of the observed network community structure, the null model was used as a benchmark to eliminate the influence of network density on the calculation of modularity [6, 7]. Specifically, we use the configuration model [25, 27] to generate a random network with the same degree sequence as the original network. Given a network, $G=(V,E)$ where $V$ is the set of nodes, $E$ is the set of edges and the degree of node $i$ is $d_{i}$. The configuration model randomly reconnects edges in the network by keeping the degree distribution of nodes constant to generate a network with the same degree sequence but randomly connected $G^{\prime}=(V,E^{\prime})$. For the original network $G$ and each generated random network $G^{\prime}_{m}(m=1,2,\dots,M)$, the same community detection algorithm (such as the Louvain algorithm [3]) is used to partition the community, and the corresponding modularity $Q$ is calculated. For the $M$ random networks $G^{\prime}_{1},G^{\prime}_{2},\dots,G^{\prime}_{M}$, calculating the modularity of the corresponding $Q^{\prime}_{1},Q^{\prime}_{2},\dots,Q^{\prime}_{M}$. Through these modularity values, we can get the distribution characteristics of the modularity of random networks, including the expected value $\mu_{Q^{\prime}}$ and the standard deviation $\sigma_{Q^{\prime}}$, where
$\mu_{Q^{\prime}}=\frac{1}{M}\sum_{m=1}^{M}Q^{\prime}_{m};$
$\sigma_{Q^{\prime}}=\sqrt{\frac{1}{M-1}\sum_{m=1}^{M}\left(Q^{\prime}_{m}-\mu Q^{\prime}\right)^{2}}.$
@@ II-B Z-score calculation
Thus, we can compare the observed modularity of the network with the modularity of this group of random networks to determine whether the network does have a strong community structure, or whether the high modularity of the network is caused by the low density of edges. We measure the significance of modularity using a Z-score. It is defined by:
$Z=\frac{Q_{\mathrm{obs}}-\mu_{Q^{\prime}}}{\sigma_{Q^{\prime}}},$
where $Q_{\mathrm{obs}}$ is the modularity of the original network.
II-C Detecting communities in temporal network
For temporal networks represented by a sequence of snapshots $G=(G_{1},G_{2},\dots,G_{t})$, we can directly apply the community detection algorithm to each time slice. In this experiment, we apply the Louvain algorithm to detect the community structure on each slice. The Louvain algorithm maximises the modularity of the partition of the node set and has $O(n\log n)$ time complexity and can detect communities quickly on large-scale networks. It is worth noting that at this step other community detection algorithms can be substituted, which does not affect the measure of community structure similarity in temporal networks that we present in this paper.
II-D NMI and the extensions
@@
Figure 1: The diagram shows the principles of UNMI and INMI respectively. In the two networks, the nodes labelled 6 and 9 appear in $G_{1}$ but are removed in $G_{2}$. The nodes labelled 1 and 4 are new in $G_{2}$. (a) UNMI compares the union of two node sets and adds a community labelled with $m^{(1)}$ in the partition of $G_{1}$, including nodes 6, and 9. A new community labelled $m^{(2)}$ has been added to the partition of $G_{2}$, including nodes 1 and 4. (B) INMI compares the intersection of two node sets, ignoring the added/removed nodes. In the example, nodes 1, 4, 6 and 9 are excluded when comparing the similarity.
@@
Recalling the definition of the NMI in a static graph context, given a network $G=(V,E)$ with community structures, let $N$ be the number of nodes in $G$. A community detection algorithm gives each node in $V$ a label. Let $L_{1}=\{1,2,\dots,r\}$ and $L_{2}=\{1,2,\dots,s\}$ be two such labellings where the actual value of the label is arbitrary (so two labellings are considered the same if they partition $V$ into the same groups even if the actual labels are different). Let $n_{r}^{(1)}$ be the number of nodes labelled $r$ by $L_{1}$ and $n_{s}^{(2)}$ be the number of nodes labelled $s$ in the second labelling $L_{2}$ and $n^{(1,2)}_{r,s}$ be the number of nodes labelled $r$ in $L_{1}$ and labelled $s$ in $L_{2}$. Then we have the probability $P(L_{1},r)$ represents the proportion of nodes labelled $r$ in $L_{1}$, $P(L_{2},s)$ represents the proportion of nodes labelled $s$ in $L_{2}$ and $P(L_{1},r,L_{2},s)$ represents the proportion of nodes labelled $r$ in $L_{1}$ and labelled $s$ in $L_{2}$:
$P(L_{1},r)=\frac{n^{(1)}_{r}}{N};$
$P(L_{2},s)=\frac{n^{(2)}_{s}}{N};$
$P(L_{1},r,L_{2},s)=\frac{n^{(1,2)}_{r,s}}{N}.$
@@
$I(L_{1};L_{2})$ $=\sum_{s}\sum_{r}P(L_{1},L_{2})\log\left(\frac{P(L_{1},L_{2})}{P(L_{1})P(L_{2})}\right)$
$=\sum_{r}\sum_{s}\frac{n^{(1,2)}_{r,s}}{N}\log\left(\frac{Nn^{(1,2)}_{r,s}}{n^{(1)}_{r}n^{(2)}_{s}}\right).$
Then the NMI is the mutual information normalised by the mean [^1] Different applications use the arithmetic mean and the geometric mean. Here we use the arithmetic average. of the entropies of the two labellings,
$\operatorname{NMI}=\frac{-2I(L_{1};L_{2})}{H(L_{1})+H(L_{2})}.$ (1)
This can be used as a measure of how similar the two labellings are with $\operatorname{NMI}=1$ meaning they are identical and $\operatorname{NMI}=0$ meaning they are completely dissimilar. Note that here, identical means not that the labels of $L_{1}$ and $L_{2}$ are the same but that two nodes have the same label in $L_{2}$ if and only if they have the same label in $L_{1}$.
II-D1 Union-NMI
Our first proposal is to align two node sets by taking the union set of two partitions. Formally, given two networks $G_{1}=(V_{1},E_{1})$ with labelling scheme $L_{1}$, $G_{2}=(V_{2},E_{2})$ with labelling scheme $L_{2}$. $n^{(1)}_{r}$ is the number of nodes labelled $r$ in $L_{1}$ and $n^{(2)}_{s}$ is the number of nodes labelled $s$ in $L_{2}$ and $n^{(1,2)}_{r,s}$ is the number of nodes labelled $r$ in $L_{1}$ and labelled $s$ in $L_{2}$. Nodes may be added or removed from the network between $G_{1}$ and $G_{2}$ thus $V_{1}$ and $V_{2}$ are two sets that are usually not equal and each may contain nodes not in the other, thus We have $V_{1}-V_{2}$ represents nodes in $V_{1}$ but not in $V_{2}$ and $V_{2}-V_{1}$ is nodes that are in $V_{2}$ but not in $V_{1}$. Let $N_{U}=|V_{1}\cup V_{2}|$ be the total number of nodes. UNMI evaluates the NMI of the partition of the union of node sets of two networks. It solves the problem of the inconsistent number of nodes in the two networks by pretending that nodes in $G_{1}$ but not in $G_{2}$ or vice versa are actually present and in some “virtual” community. We pretend these nodes are present in a community labelled $m^{(1)}$ in the labelling scheme $L_{1}$ and a community labelled $m^{(2)}$ in the labelling scheme $L_{2}$ (Fig 1(a)). The mutual information between the two labelling schemes is

### v3

@@ Abstract
When we detect communities in temporal networks it is important to ask questions about how they change over time. Adjusted mutual information (AMI) has been used to measure the similarity of communities when the nodes on a network do not change. We propose two extensions, namely, Union-Adjusted Mutual Information (UAMI) and Intersection-Adjusted Mutual Information (IAMI). UAMI and IAMI evaluate the similarities of community structures when nodes are added or removed. Experiments show that these methods are effective in dealing with temporal networks with the changes in the set of nodes, and can capture the dynamic evolution of community structure in both synthetic and real temporal networks. This study not only provides a new similarity measurement method for network analysis but also deepens the understanding of community change in complex temporal networks.
keywords:
Temporal networks, Dynamic communities, Network statistics
1 Introduction
@@
An important problem in network science is understanding how communities form in networks. Communities within these networks do not remain static: they expand, merge, or dissolve, reflecting the underlying changes in relationships and interactions [tardelli2024temporal, greene2010Tracking, chen2010detecting, du2015tracking]. We might hypothesise that some networks have relatively stable and long-lasting relationships but other networks might form only transient communities, but how can this be measured? A community here is a partition of a network into (usually) distinct sets of nodes with connections more common within communities. A common measure used on such a partition is modularity [newman2003structure, newman2006modularity, fortunato2007resolution], though it has attracted some criticism [peixoto2023descriptive]. When we consider a network that changes in time we can choose to analyse the communities only within a time window and ask how much communities change between such windows. This requires us to rigorously compare two different partitions of a network and to give a measure of how similar those partitions are. This measure must allow for the fact that nodes can leave or join the network.
@@
A well-known method for comparing two such partitions is adjusted mutual information (AMI) but this measure assumes that both networks have the same set of nodes. The primary challenge addressed in this research is the accurate comparison of community structures in networks where the set of nodes varies over time. This study proposes variants of AMI: UAMI and IAMI which compares two networks with different node sets either by considering the union or intersection of the node sets. We show by using artificial models that these measure different aspects of how communities change over time and demonstrate their utility on artificial networks with known characteristics. Following this we use real temporal network data sets which we split into different time windows. We demonstrate how graph statistics can be used to pick an appropriate time window for the task. The methods accurately recover the known properties of artificial data. On real-world datasets, however, community stability is generally brief—shorter than in our synthetic models—and varies across different cases.
2 Related work
@@
The focus of this paper is on how communities change rather than specifically on how to partition a network into communities. Here we begin with a basic assumption, commonly made, that a community detection algorithm partitions all nodes in a network into a set of non-overlapping communities. Community detection is far from a solved problem. The problem can be divided into inferential methods and descriptive methods [peixoto2023descriptive]. Inferential methods attempt to fit parameters in some predetermined model that generates communities. Descriptive methods attempt to partition the network to maximise some measure of how good that partition is at creating communities. Modularity [newman2006modularity] is commonly used for the latter as it attempts to measure the proportion of links that start and end in the same community rather than that which would be expected by a null model. However, modularity suffers from a resolution limit [fortunato2007resolution], causing small but meaningful communities to be merged in large networks. To mitigate this problem, previous research introduced modularity density to better capture within-community link compactness, and adopted z-score to evaluates the statistical significance of modularity scores against a configuration-model ensemble [chen2014community, atsushi2016zscore]. In this paper, the measures we use do not rely on whether descriptive or inferential techniques are available. We use the Louvain algorithm to detect communities for our investigation into real-world data because it is well-known, relatively quick to run, and simple to understand. However, the core contribution of the UAMI and IAMI and their use to measure how fast communities change in time would not change if a different community detection method were used [raghavan2007near, rosvall2008maps]. For the work on artificial networks, we simply assume that the communities exist (community membership is generated directly from the model) and that some algorithm could in principle detect them.
The dynamic characteristics of dynamic communities play an important role in the study of the evolution of complex systems, which provides a bridge for the interaction between micro and macro structures in complex systems. To effectively capture these evolving structures over time, researchers commonly use the snapshot model to analyse dynamic networks. Each snapshot corresponds to a static network, and the community structure on each snapshot is obtained by using the static community detection algorithm, then the evolution of the community structure over time can be analysed by comparing the structure on different snapshot. Thus, the choice of the size of time window is important. A small time window can capture short-term fluctuations [krings2012effects], but it is subject to noise interference. A large time window can smooth out short-term fluctuations, but it will lose local details [cleveland2013smoothing]. The paper [peel2015detecting] applies the changing point detection method to locate the moments when the network structure undergoes significant changes, providing a data-driven basis for the division of the time window. The paper [cazabet2021data] utilises the principle of data compression and selects the dynamic network expression form that achieves the best compromise between representation efficiency and interpretability. Paper [peixoto2017modelling] proposed a model based on dynamic stochastic block model that can non-parametically infer the number of communities and their evolution time scale in a continuous-time or non-stationary background, without the need to set the window length in advance. The paper [delvenne2010stability] discusses the stability of the community structure in networks and proposes a stability measurement method based on the Markov process. This method reveals the community structure characteristics of the network at different resolutions through time-scale analysis.
@@
Based on existing studies, we improved existing techniques and compared the similarity of community structure of temporal networks at different time points on a one-to-one basis. A direct approach to comparing the similarity of networks at the mesoscale level is comparing the similarity of the partition of nodes. Normalised Mutual Information (NMI) [danon2005comparing] is a widely used method to measure the similarity of partitions, by comparing the similarity of two partitions by measuring information gain. Recent studies have pointed out that the NMI is biased, and its variants [vinh2009information, amelio2015normalized] and asymmetric normalise technique [jerdee2023normalized] are proposed to solve the problem. In the context of overlapping communities, where nodes can have more than one community labels, Overlapping NMI can be used [mcdaid2011normalized, lancichinetti2009detecting] to measure the similarity in community structures. In addition, partition edit distance (PED) [aynaud2010static] is an extension of graph edit distance that defines operands to manipulate the partition of nodes. However, it relies on one-to-one matching between communities, and the high computational overhead also limits the application of this method on large-scale networks.
3 Methods
@@
Comparing network partitions is a fundamental task in network science and community detection. Mutual Information (MI) and its extensions are widely used metrics due to their ability to effectively quantify the similarity between partitions. However, mutual information assumes that the partitions being compared share an identical set of nodes, which often is not the case in real-world scenarios where networks may have differing node sets due to growth, decay, or entirely different origins. This limitation becomes evident when attempting to compare partitions of networks with varying node compositions, rendering traditional mutual information inadequate for such analyses. Recognising this gap, we propose an approach that explicitly compares network partitions with different node sets by introducing two novel metrics: Union-AMI (UAMI) and Intersection-AMI (IAMI). This approach is motivated by the practical need to assess the similarity between partitions in networks where node addition or deletion is common, such as in temporal networks or comparative studies across different systems.
@@
In our framework, UAMI considers the union of the node sets from the partitions being compared, effectively incorporating all nodes present in either partition. This allows for a comprehensive evaluation of similarity that accounts for the entirety of both networks. Conversely, IAMI focuses on the intersection of the node sets, assessing similarity based solely on the nodes common to both partitions.
3.1 Temporal networks model
@@
We adopt the snapshot model [holme2019temporal] of a sequential network to study the dynamic evolution of networks in time dimension. The model works by dividing the entire temporal network into a series of discrete time segments, each corresponding to a snapshot of the temporal network. This method allows us to observe changes in the structure of the network at different points in time and analyse the dynamic behaviour of nodes and edges. Here we consider our graph to be constructed from a series of events (edges) that connect two nodes and have a given time of occurrence. A temporal network that is observed between $t_{0}$ and $t_{\max}$ can be represented by a series of snapshots $(G_{1},G_{2},\dots,G_{n})$ where each snapshot contains only the edges in the time window $T_{i}\in[t,t+\tau)$, denoted by , where $\tau$ is the size of the time window. Each snapshot $G_{i}=(V_{i},E_{i})$ represents the subgraph in time window $T_{i}$. $V_{i}$ are the nodes $E_{i}$ are the edges are active in time window $T_{i}$. We can also modify this definition slightly to allow snapshots where the time windows overlap.
Choosing the right window size is critical to building a snapshot model of a dynamic network. A small time window will result in sparse graphs within snapshots as the number of edges within the window will be small. A large time window smooths out short-term fluctuations, thereby losing local information. Although several methods for extracting timescales in dynamic data sets exist already [darst2016detection, peel2015detecting, cazabet2021data, peixoto2017modelling], given that our problem is choosing a reasonable time window to build a network with meaningful community structure, we analyse the modularity of the network under different windows and the proportion of the largest connected component to select the appropriate time window. Intuitively, in a snapshot model constructed with a reasonable time window, the modularity of each time slice should be large, indicating that there is a meaningful community structure in the network at this time. The research [fortunato2007resolution] indicates that modularity is not fair when comparing networks of different sizes and densities, and this is known as the resolution limit of modularity. Therefore, we eliminate the bias in modularity caused by networks of different sizes and densities by comparing the z-scores [chen2015new, chen2018network, atsushi2016zscore] obtained from the null model [newman2003structure]. Given a network, $G=(V,E)$ where $V$ is the set of nodes, $E$ is the set of edges and the degree of node $i$ is $d_{i}$. The configuration model randomly reconnects edges in the network by keeping the degree distribution of nodes constant to generate a network with the same degree sequence but randomly connected $G^{\prime}=(V,E^{\prime})$. For the original network $G$ and each generated random network $G^{\prime}_{m}(m=1,2,\dots,M)$, the same community detection algorithm is used to partition the community, and the corresponding modularity $Q$ is calculated. For the $M$ random networks $G^{\prime}_{1},G^{\prime}_{2},\dots,G^{\prime}_{M}$, calculating the modularity of the corresponding $Q^{\prime}_{1},Q^{\prime}_{2},\dots,Q^{\prime}_{M}$. Through these modularity values, we can get the distribution characteristics of the modularity of random networks, including the expected value $\mu_{Q^{\prime}}$ and the standard deviation $\sigma_{Q^{\prime}}$, where
$\mu_{Q^{\prime}}=\frac{1}{M}\sum_{m=1}^{M}Q^{\prime}_{m};\quad\sigma_{Q^{\prime}}=\sqrt{\frac{1}{M-1}\sum_{m=1}^{M}\left(Q^{\prime}_{m}-\mu Q^{\prime}\right)^{2}}.$
@@
Thus, we can compare the observed modularity of the network with the modularity of this group of random networks to determine whether the network does have a strong community structure, or whether the high modularity of the network is caused by the low density of edges. We measure the significance of modularity using a z-score. It is defined by: $Z=\frac{Q_{\mathrm{obs}}-\mu_{Q^{\prime}}}{\sigma_{Q^{\prime}}}$, where $Q_{\mathrm{obs}}$ is the modularity of the original network.
3.2 Mutual information and the extensions
@@
Figure 1: The diagram shows the principles of union extension and intersection extension respectively. In the two networks, the nodes labelled 6 and 9 appear in $G_{1}$ but are removed in $G_{2}$. The nodes labelled 1 and 4 are new in $G_{2}$. (a) The union extension compares the union of two node sets and adds a community labelled with $m^{(1)}$ in the partition of $G_{1}$, including nodes 6, and 9. A new community labelled $m^{(2)}$ is added to the partition of $G_{2}$, including nodes 1 and 4. (B) The intersection extension compares the intersection of two node sets, ignoring the added/removed nodes. In the example, nodes 1, 4, 6 and 9 are excluded when comparing the similarity.
@@
Recalling the definition of the mutual information in a static graph context, given a network $G=(V,E)$ with community structures, let $N$ be the number of nodes in $G$. A community detection algorithm gives each node in $V$ a label. Let $L_{1}=\{1,2,\dots,r\}$ and $L_{2}=\{1,2,\dots,s\}$ be two such labellings where the actual value of the label is arbitrary (so two labellings are considered the same if they partition $V$ into the same groups even if the actual labels are different). Let $n_{r}^{(1)}$ be the number of nodes labelled $r$ by $L_{1}$ and $n_{s}^{(2)}$ be the number of nodes labelled $s$ in the second labelling $L_{2}$ and $n^{(1,2)}_{r,s}$ be the number of nodes labelled $r$ in $L_{1}$ and labelled $s$ in $L_{2}$. Then we have the probability $P(L_{1},r)$ represents the proportion of nodes labelled $r$ in $L_{1}$, $P(L_{2},s)$ represents the proportion of nodes labelled $s$ in $L_{2}$ and $P(L_{1},r,L_{2},s)$ represents the proportion of nodes labelled $r$ in $L_{1}$ and labelled $s$ in $L_{2}$:
$P(L_{1},r)=\frac{n^{(1)}_{r}}{N};P(L_{2},s)=\frac{n^{(2)}_{s}}{N};P(L_{1},r,L_{2},s)=\frac{n^{(1,2)}_{r,s}}{N}.$
@@
$I(L_{1};L_{2})=\sum_{s}\sum_{r}P(L_{1},L_{2})\log\left(\frac{P(L_{1},L_{2})}{P(L_{1})P(L_{2})}\right)=\sum_{r}\sum_{s}\frac{n^{(1,2)}_{r,s}}{N}\log\left(\frac{Nn^{(1,2)}_{r,s}}{n^{(1)}_{r}n^{(2)}_{s}}\right)$.
One extension of mutual information is the Normalised Mutual Information (NMI), which is obtained by normalising the mutual information by the mean of the entropies of the two labelings. Depending on the application, either the arithmetic mean or the geometric mean may be used for normalisation. In this work, we adopt the arithmetic mean. Recent studies have shown that NMI can be biased by the number of communities, which means that NMI tends to give high scores to communities with a large number of small partitions [vinh2009information, jerdee2023normalized, mccarthy2019metrics]. Thus, adjusted mutual information (AMI) [vinh2009information] is proposed to correct for this bias, pointed out that the bias can be eliminated by subtracting the expected value under the null model configuration, the expectation term can be expressed by:
$\mathbb{E}[\operatorname{MI}]=\sum_{r=L_{1}}\sum_{s=L_{2}}\sum_{n=n_{r,s}^{\min}}^{n_{r,s}^{\max}}\frac{\binom{n_{r}^{(1)}}{n}\binom{N-n_{r}^{(1)}}{n_{s}^{(2)}-n}}{\binom{N}{n_{s}^{(2)}}}\times\frac{n}{N}\log\left(\frac{nN}{n_{r}^{(1)}n_{s}^{(2)}}\right).$ (1)
Then the expression of AMI is $\frac{2(\operatorname{MI}-\mathbb{E}[\operatorname{MI}])}{H(L_{1})+H(L_{2})-\mathbb{E}[\operatorname{MI}]}$. The AMI takes a value of 1 when the two partitions are identical and 0 when the MI between two partitions equals the value expected due to chance alone. In the following expansion, we will apply the same operation to eliminate the bias caused by the number of communities.
3.2.1 Union-AMI
Our first proposal is to align two node sets by taking the union set of two partitions. Formally, given two networks $G_{1}=(V_{1},E_{1})$ with labelling scheme $L_{1}$, $G_{2}=(V_{2},E_{2})$ with labelling scheme $L_{2}$. $n^{(1)}_{r}$ is the number of nodes labelled $r$ in $L_{1}$ and $n^{(2)}_{s}$ is the number of nodes labelled $s$ in $L_{2}$ and $n^{(1,2)}_{r,s}$ is the number of nodes labelled $r$ in $L_{1}$ and labelled $s$ in $L_{2}$. Nodes may be added or removed from the network between $G_{1}$ and $G_{2}$ thus $V_{1}$ and $V_{2}$ are two sets that are usually not equal and each may contain nodes not in the other, thus We have $V_{1}-V_{2}$ represents nodes in $V_{1}$ but not in $V_{2}$ and $V_{2}-V_{1}$ is nodes that are in $V_{2}$ but not in $V_{1}$. Let $N_{U}=|V_{1}\cup V_{2}|$ be the total number of nodes. The union extension solves the problem of the inconsistent number of nodes in the two networks by pretending that nodes in $G_{1}$ but not in $G_{2}$ or vice versa are actually present and in some “virtual" community. We pretend these nodes are present in a community labelled $m^{(1)}$ in the labelling scheme $L_{1}$ and a community labelled $m^{(2)}$ in the labelling scheme $L_{2}$ (Fig 1(a)). The mutual information between the two labelling schemes is